1) Find the area bounded by the curve y=9-x and the x axis.




a) Horizontal Strip





b) Vertical Strip

1
Expert's answer
2022-05-26T07:56:24-0400

1)



"9-x^2=0=>x_1=-3, x_2=3"

a)



"y=9-x^2=>x=-\sqrt{9-y}\ or \ x=\sqrt{9-y}""A=\displaystyle\int_{0}^{9}(\sqrt{9-y}-(-\sqrt{9-y}))dy""=2\displaystyle\int_{0}^{9}\sqrt{9-y}dy=2\big[-\dfrac{2}{3}(9-y)^{3/2}\big]\begin{matrix}
 9 \\
 0
\end{matrix}""=-\dfrac{4}{3}((9-9)^{3/2}-(9-0)^{3/2})=36({units}^2)"


b)



"A=\displaystyle\int_{-3}^{3}(9-x^2)dx=\big[9x-\dfrac{x^3}{3}\big]\begin{matrix}
 3 \\
 -3
\end{matrix}""9(3)-\dfrac{(3)^3}{3}-(9(-3)-\dfrac{(-3)^3}{3})=36({units}^2)"